hello i just wanted to check to see if I was doing this word problem correctly. The word problem says...


A tank can be filled by an inlet pipe in 8 hours when the drain pipe is open and in 5 hours if the drain pipe is closed. How long will it take to empty a full tank when the inlet pipe is closed.

i did...

X/8 (x over eight) + X/5 (x over five) = 1

40 [X/8 + X/5 = 1]

then got...

5x + 8x = 40
13x= 40
x= 3 1/3


if this is correct, wrong, or an easier way to do thi problem it would be very helpful to know. Thank You!

Answer by Edwin McCravy(20054)


A tank can be filled by an inlet pipe in 8 hours when the drain pipe 
is open and in 5 hours if the drain pipe is closed. How long will it 
take to empty a full tank when the inlet pipe is closed. 

I'm afraid your method is wrong.  Only 3 1/3 hours is much too fast 
for the drain pipe to be able to empty the pool if it only slows 
the fill pipe down from filling it in 5 hours to filling it in 8 
hours. If it drained that fast, the inlet pipe would never be able 
to fill the pool at all when the drain pipe was open for it would
be draining water out much faster than the inlet pipe was putting 
water in!

Here's a method very much like distance-rate-time problems that I 
think you may understand better.

Let's make this DRT-chart, except here the D stands for "Deeds Done" 
rather than "Distance".  So D represents the number of deeds, or 
poolfuls it caused.

                      D             R           T  
                    No. of       Rate in    No. of hrs.  
                   poolfuls      pool/hr    required
Inlet pipe alone                                  
Drain pipe alone                                
both open                                       

Let x = the length of time required for the drain pipe to empty the
pool when the inlet pipe is open.

First fill in the three times, 5 hours for the inlet pipe, x hours
for the drain pipe, and 8 hours when they are both open.

                      D             R           T
                    No. of       Rate in    No. of hrs.  
                   poolfuls      pool/hr    required
Inlet pipe alone                               5
Drain pipe alone                               x
both open                                      8

Next we will fill in the number of poolfuls.

The inlet pipe caused the GAIN of 1 poolful in 5 hours.  So we put
1 for the number of poolfuls it caused in the 5 hours it was open
by itself.

                      D             R           T
                    No. of       Rate in    No. of hrs.  
                   poolfuls      pool/hr    required
Inlet pipe alone        1                      5
Drain pipe alone                               x
both open                                      8

Now the drain pipe caused a LOSS of 1 poolful in x hours. So we
put -1 for the number of poolfuls it caused in the x hours it was
open by itself.  (That -1 is the tricky part!)

                      D             R           T
                    No. of       Rate in    No. of hrs.  
                   poolfuls      pool/hr    required
Inlet pipe alone        1                      5
Drain pipe alone       -1                      x
both open                                      8

Now when both were open that caused a GAIN of 1 poolful in 8 hours.
So we put 1 for the number of poolfuls they caused in the 8 hours 
they both were open. So we fill in 1 for the no. of poolfuls caused 
by both of them being open.

                      D             R           T
                    No. of       Rate in    No. of hrs.  
                   poolfuls      pool/hr    required
Inlet pipe alone        1                      5
Drain pipe alone       -1                      x
both open               1                      8

Now we fill in the rates by using R = D/T.

                      D             R           T
                    No. of       Rate in    No. of hrs.  
                   poolfuls      pool/hr    required
Inlet pipe alone        1          1/5         5
Drain pipe alone       -1         -1/x         x
both open               1          1/8         8

Now the logic in setting up the equation is just like
when a boat is in a river going downstream, we add the
rate of the boat in still water and the rate of the 
stream to get the combined rate.  So we use the same 
idea here:

Rate of inlet pipe + Rate of drain pipe = their combined rate, or

             1/5   +        (-1/x)      =        1/8

                              1/5 - 1/x = 1/8 

Solve that and you'll get 13 1/3 hours for the drain pipe to 
drain 1 full pool.  That answer makes a lot more sense.

Edwin